One reaction that produces hydrogen gas can
be represented by the unbalanced chemical
equation
Mg(s) + HCl(aq) −! MgCl2(aq) + H2(g) .
What mass of HCl is consumed by the reac-
tion of 3.17 mol of magnesium?
Answer in units of g
1
Expert's answer
2012-12-07T04:11:40-0500
Find the amount of the substance magnesium:
n(Mg) = m/M
n(Mg) = 3.17/24 = 0.132083333 mol
Mg(s) + 2HCl(aq) −! MgCl2(aq) + H2(g)
n(Mg) = 2n(HCl) = 0.2641mol
m(HCl) = M*n
m(HCl) = 36.5*0.2641 = 9.63965 g